OpenStudy (anonymous):

Find the quotient and simplify: X+3/ x^2 ÷ (x^2-9)/ 4x^2

OpenStudy (jackthegreatest):

hi ashley i can help

OpenStudy (anonymous):

Hey! Thank you!

OpenStudy (jackthegreatest):

so right now, do u have any idea wat ur first step should be?

OpenStudy (anonymous):

Im really not sure

OpenStudy (jackthegreatest):

k

OpenStudy (jackthegreatest):

well first of all, u dont want a fraction divided by a fraction

OpenStudy (anonymous):

Alright

OpenStudy (jackthegreatest):

so u could get rid of the second fraction by flipping the top with the bottom

OpenStudy (anonymous):

Will the denominator be x-3 ?

OpenStudy (jackthegreatest):

\[\frac{ x+3 }{ x^{2} } \times \frac{ x^{2} - 9 }{ 4x^{2} }\]

OpenStudy (jackthegreatest):

do u know how to cross multiply?

OpenStudy (anonymous):

Yes

OpenStudy (jackthegreatest):

actually after thinking about it, u dont need to cross multiply, just multiply the top with the top and the bottom with the bottom

OpenStudy (anonymous):

Im not sure im following

OpenStudy (jackthegreatest):

Hold on a sec

OpenStudy (jackthegreatest):

I'm gonna post a pic of my steps

OpenStudy (anonymous):

Okayy

OpenStudy (anonymous):

What would the denominator be?

OpenStudy (jackthegreatest):

there finally i got it by using mail from my phone to computer

OpenStudy (anonymous):

I understand thay now that you for helping me !!

OpenStudy (jackthegreatest):

welcome!

OpenStudy (anonymous):

Okay i have another problem. X/2 ÷ 5y/2 would the answer be x/5y ??

OpenStudy (anonymous):

@jackthegreatest

OpenStudy (jackthegreatest):

yep!! congrats! u got it right!

OpenStudy (anonymous):

Yay thank you!

OpenStudy (jackthegreatest):

if any more questions come up just post up a new questions, that way more people could help u as well : )

OpenStudy (jackthegreatest):

and just tag me and i shall be summoned... lol

OpenStudy (anonymous):

Lol okay thank you again